Using superposition and available deflection table determine the rotation at A and deflection at C.

4 KN 4 KN dr-
Answer #1

* Ansi - super position : ♡ on slope at 8) (11) Ye (deflection at ) . UkN UKN DB kt r 4m um By . UKN YKN с Ова t- Iyz om- 27from table, OAL = ; A = 4x6 (82-6?) Pb (2_62) Pb (2_6?); PA, = 4* BFI (8)2 6 EIL 131 acnice Boy cncl x=4n 6 ETI primer cop emUKN Comm= Pb (26) - GEIL 842 = 472[9-23 - 10 (6*8) EI 7539 Yeni= 4 x 244) (8² - 4 - 2 S 6EI () : = 2407 iyo = to om + 5 Jebo